1. Kinkaid Co. is incorporated at the beginning of this year and engages in a number of transactions. The following journal entries impacted its stockholders’ equity during its first year of operations.

General Journal Debit Credit
a. Cash 270,000
Common Stock, $25 Par Value 240,000
Paid-In Capital in Excess of Par Value, Common Stock 30,000
b. Organization Expenses 180,000
Common Stock, $25 Par Value 127,000
Paid-In Capital in Excess of Par Value, Common Stock 53,000
c. Cash 43,500
Accounts Receivable 17,000
Building 82,800
Notes Payable 59,500
Common Stock, $25 Par Value 53,800
Paid-In Capital in Excess of Par Value, Common Stock 30,000
d. Cash 120,000
Common Stock, $25 Par Value 78,000
Paid-In Capital in Excess of Par Value, Common Stock 42,000


Required:
2. How many shares of common stock are outstanding at year-end?
3. What is the amount of minimum legal capital (based on par value) at year-end?
4. What is the total paid-in capital at year-end?
5. What is the book value per share of the common stock at year-end if total paid-in capital plus retained earnings equals $791,000?

2.

Kohler Corporation reports the following components of stockholders’ equity on December 31, 2016:

Common stock—$15 par value, 100,000 shares authorized,
45,000 shares issued and outstanding
$675,000
Paid-in capital in excess of par value, common stock 60,000
Retained earnings 430,000
Total stockholders' equity $1,165,000


In year 2017, the following transactions affected its stockholders’ equity accounts.

Jan. 1 Purchased 5,500 shares of its own stock at $15 cash per share.
Jan. 5 Directors declared a $4 per share cash dividend payable on February 28 to the February 5 stockholders of record.
Feb. 28 Paid the dividend declared on January 5.
July 6 Sold 2,063 of its treasury shares at $19 cash per share.
Aug. 22 Sold 3,437 of its treasury shares at $12 cash per share.
Sept. 5 Directors declared a $4 per share cash dividend payable on October 28 to the September 25 stockholders of record.
Oct. 28 Paid the dividend declared on September 5.
Dec. 31 Closed the $428,000 credit balance (from net income) in the Income Summary account to Retained Earnings.


Required:

1. Prepare journal entries to record each of these transactions for 2017.
2. Prepare a statement of retained earnings for the year ended December 31, 2017.
3. Prepare the stockholders' equity section of the company’s balance sheet as of December 31, 2017.
Answer #1

1a) Share outstanding = (240000+127000+53800+78000)/25 = 19952 Shares

1b) Minimum legal paid in capital = 19952*25 = $498800

1c) Total paid in capital = 498800+30000+53000+30000+42000 = $653800

1d) Book value per share = (653800+791000)/19952 = 72.41 per share

Note : Please post each question individually as per HomeworkLib guidelines